OpenStudy (anonymous):

Find the GCF of -10c^2 d and 15cd^2. Any help?

OpenStudy (anonymous):

factors of -10 =-2,5 factors of 15=3*5 common to both=3 c^2d=c*c*d cd^2=c*d*d common to these =c*d Hence GCF=?

OpenStudy (anonymous):

common write common to both=5 not 3